OpenStudy (anonymous):

Which of the following is equivalent to five and four-ninths divided by seven?

OpenStudy (anonymous):

what are the following??

OpenStudy (anonymous):

seven-ninths one and two-sevenths five-sevenths five and four sixty-thirds

OpenStudy (anonymous):

\[\huge \frac{5 \frac{4}{9} } {7} =\frac{49}{9 \times 7} = \frac{7}{9}\]

OpenStudy (anonymous):

Hence seven-ninths is the correct option..

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

thank yall

OpenStudy (anonymous):

np.. ( no problem)